solve and check 2(x-4)+3(x+5)=2x-2
Answer by blwinbbbles(106)

solve:
2(x-4) + 3(x+5) = 2x-2 so first thing is simplify the equation
2x-8 + 3x+15 = 2x -2 keep simplifying
5x + 7 = 2x - 2 subtract 2x from both sides
3x + 7 = -2 subtract 7 from both sides
3x = -9 divide by 3
x = -3
Check:
2(-3-4) + 3(-3+5) =2(-3)-2
2(-7) + 3(2) = -6-2
-14 + 6 = -8
-8 = -8
so we know that the answer x = -3 is correct